# Adaptive Canonicalization With Application to Invariant Anisotropic Geometric Networks

### v1

@@ B.1 Universal Approximation of Euclidean Fucntions
Theorem 14 (Universal Approximation Theorem [Leshno et al., 1993]).
@@ B.2 Universal Approximation of Multi-Sets Functions
For example, DeeptSets are universal approximators of invariant $C_{0}(\mathcal{K},{\mathbb{R}}^{D})$ functions [Wagstaff et al., 2022].
@@ Appendix D Functional Calculus and Spectral Filters
applies the graph convolution theorem [Bracewell and Kahn, 1966] with $d^{\prime}$ output channel: each Fourier mode $v_{i}$ is preserved in space, while channels are mixed by $g(\lambda_{i})$ in the spectral domain. In the scalar case ($T=d^{\prime}=1$) with $f:\mathbb{R}\to\mathbb{R}$, the spectral filter simply reduces to the functional-calculus operator acting on $\mathbf{X}$:
@@
Spectral graph neural networks [Defferrard et al., 2016b, Kipf, 2016, Levie et al., 2018] compose such filters with pointwise nonlinearities, using trainable $g$ at each layer.
Automorphism Invariance of Spectral Filters.
By construction, a spectral filter depends only on the eigenvalues and eigenvectors of the GSO $\mathcal{L}$. Let $\pi$ be a graph automorphism (i.e., a permutation of vertices that preserves adjacency and hence commutes with $\mathcal{L}$), the operator satisfies
$P_{\pi}\mathcal{L}P_{\pi}^{\top}=\mathcal{L},$
where $P_{\pi}$ is the permutation matrix associated with $\pi$. By functional calculus, $f(\mathcal{L})=P_{\pi}f(\mathcal{L})P_{\pi}^{\top}$ holds for any spectral filter $g$. This means that the action of a spectral filter does not change under any automorphism of the graph, i.e., its output is invariant to node re-labeling.
When we apply a spectral filter on a square grid graph, the response along the $x$-axis and along the $y$-axis is the same. This implies that the spectral filters are isotropic in the spatial domain, and therefore, the spectral filters can not distinguish between directions, such as horizontal vs. vertical, and impose directional limitations.
Appendix E Application of Adaptive Canonicalization to Anisotropic Geometric Networks: Construction Details
@@
In this section, we present the construction details of the application of adaptive canonicalization to anisotropic geometric networks
E.1 Construction Details for Anisotropic Nonlinear Spectral Filters
We introduce anisotropic nonlinear spectral filters as a mechanism to perform adaptive canonicalization for eigenbasis selection within the framework of spectral filtering. Unlike standard approaches, which apply uniform (isotropic) filtering across the spectrum, our formulation partitions the spectrum into disjoint frequency bands and treats each independently. Within every band, we introduce a unitary transformation that reorients the eigenspace based on prior maximization, thereby resolving orientation ambiguities inherent in spectral methods.
In the spectral filter, we consider the GSO as the normalized graph Laplacian $\mathcal{L}=\mathbf{I}-\mathbf{D}^{-\frac{1}{2}}\mathbf{A}\mathbf{D}^{-\frac{1}{2}}$, where $\mathbf{A}$ is the adjacency matrix and $\mathbf{D}$ is a diagonal degree matrix with entries ${\mathbf{d}}=(d_{i})_{i}\in\mathbb{R}^{N}$. The eigendecomposition of $\mathcal{L}$ is given by $\mathcal{L}={\mathbf{V}}\mathbf{\Lambda}{\mathbf{V}}^{\top}$, where ${\mathbf{V}}=({\mathbf{v}}_{i})_{i}\in{\mathbb{R}}^{N\times N}$ is the matrix of eigenvectors (i.e., the eigenbasis) and $\mathbf{\Lambda}=\operatorname{diag}(\lambda_{1},\ldots,\lambda_{N})$ is the diagonal matrix of eigenvalues, where $\lambda_{1}\leq\ldots\leq\lambda_{N}$. Consider predefined bands $b_{0}<b_{1}<\ldots<b_{B}$ where $b_{0},b_{B}]\subseteq[0,2]$. In particular, we adopt a dyadic partitioning scheme where the subbands are defined as $\left[\lambda_{N}r^{S-k+1},\lambda_{N}r^{S-k}\right]$ with $0<r<1$ denoting the decay rate, following the approach introduced in Lin et al. [2024b]. We note that band design can be realized through different mechanisms, such as an attention mechanism, e.g., SpecFormer [Bo et al., 2023]. We leave this exploration for future work.
The indicator functions $P_{k}:=\mathbbm{1}_{[b_{k-1},b_{k})}:{\mathbb{R}}\rightarrow{\mathbb{R}}_{+}$ are the orthogonal projections onto the eigenspace corresponding to the $k$-th band. We augment these projections with band-wise unitary actions that commute with the GSO $\mathcal{L}$. Specifically, for each spectral band $k$ we consider a unitary matrix ${\mathbf{U}}_{k}\in\mathcal{O}(M_{k})$, where $M_{k}=\mathrm{rank}(P_{k})$. The global transformation operator is thus block-diagonal in the spectral domain ${\mathbf{U}}=\sum_{k}{\mathbf{U}}_{k}P_{k}\equiv\bigoplus_{j}{\mathbf{U}}_{k}$. Each eigenspace is treated independently but with a learnable, band-specific orientation.
We consider a classification task with $D$ classes and deploy the prior maximization as follows. Specifically, we introduce $D$ class-specific priors and orientations. For each class $d\in\{1,\ldots,D\}$ and band $k$, let $\mathbf{U}^{(d)}_{k}\in\mathcal{O}(M_{k})$ and define the block operator $\mathbf{U}^{(d)}\!\!=\bigoplus_{k=1}^{B}\mathbf{U}^{(d)}_{k}$. Let ${\mathbf{X}}\in{\mathbb{R}}^{N\times T}$ be the input with $T$ channels per node and let ${\mathbf{V}}_{k}\in\mathbb{R}^{N\times M_{k}}$ be the matrix of eigenvectors spanning the $k$-th band. The class-$d$ band coefficients are
$\widetilde{\mathbf{X}}^{(d)}_{k}=\mathbf{U}^{(d)}_{k}\,\mathbf{V}_{k}^{\top}\mathbf{X}\in\mathbb{R}^{M_{k}\times T}.$
Each class has its own prior that selects an orientation that stresses features for that class. We denote $h_{d}$ the class-$d$ prior. The class-$d$ orientation is obtained by
$\mathbf{U}^{\square(d)}=\bigoplus_{k}{\mathbf{U}}_{k}^{\square(d)}=\underset{\mathbf{U}^{(d)}_{1},\ldots,\mathbf{U}^{(d)}_{B}}{\arg\max}\;h_{d}(\Psi_{d}(\phi(\widetilde{\mathbf{X}}^{(d)}_{1},\ldots,\widetilde{\mathbf{X}}^{(d)}_{B})))$,
where $\phi$ is a mapping on band-wise features, and $\Psi_{d}$ is the class-specific classifier for class $d$. The nonlinear aspect emerges from the mapping network on the spectral coefficients, which operates on the band-wise features. The optimal orientations are then applied band-wise and class-wise, and the network output is computed with respect to this canonical eigenbasis. The anisotropy stems from the independent reorientation of spectral bands, in contrast to isotropic formulations that enforce a single global alignment of the spectrum.
One vs. Rest (Independent Binary Classifiers).
In each class-specific classifier, we apply a scoring function to the prior’s output. Specifically, we consider the one vs. rest setup [Rifkin and Klautau, 2004, Galar et al., 2011, Allwein et al., 2000]. We obtain the class logit
$s_{d}=\Psi_{d}(\phi(\widetilde{\mathbf{X}}^{\square(d)}_{1},\ldots,\widetilde{\mathbf{X}}^{\square(d)}_{B})),\quad\widetilde{\mathbf{X}}^{\square(d)}_{k}:=\mathbf{U}^{\square(d)}_{k}\,\mathbf{V}_{k}^{\top}\mathbf{X}.$
@@ E.1 Construction Details for Anisotropic Nonlinear Spectral Filters
We then define the per-class score by a sigmoid $\hat{y}_{d}=\sigma(s_{d})$. Note that the vector $(\hat{y}_{1},\ldots,\hat{y}_{D})$ is not a probability measure, since each entry represents the independent probability of class $c$ being present. We use binary cross-entropy per class and sum over classes:
$\sum_{d=1}^{D}(-\,y_{d}\log\sigma(s_{d})-(1-y_{d})\log(1-\sigma(s_{d}))).$
Note that we can include class weights $w_{d}\geq 0$ to handle imbalance between classes. Then, at test time we predict the class with the largest logit $\arg\max_{d}s_{d}$.
We develop the graph-level representation of adaptive canonicalization applied to spectral methods, which is fully spectral. Details on the corresponding node-level representation are provided in App. G.2.
Optimizing Unitary Orientations for Band-Specific Representation.
For multiple graphs, $M_{k}$ may vary; to standardize dimensions, each MLP $\phi$ uses the largest band size and smaller bands are zero-padded. From an implementation standpoint, we can search over unitary orientations $\{{\mathbf{U}}_{k}^{(\cdot)}\}_{k=1}^{B}$, one per band, and select the pooled set of orientations that maximizes the prior coupled with the classification task. One option is to generate a set of Haar-distributed random unitary matrices for each band [Mezzadri, 2006]. Based on the chosen prior objective, the network then selects the orientation to retain. This procedure, sampling random unitaries and picking the one that maximizes the prior, should be viewed as one possible implementation rather than a requirement. Alternatively, the prior can be maximized more effectively by refining the sampled matrices. For example, we can initialize with random unitary matrices, apply the prior maximization, and then perform a few additional gradient descent (GD) steps to reduce the objective locally. In this case, the optimization is treated as an inner loop: the refined orthogonal matrix is plugged into the outer computation, without differentiating through the iterative gradient descent itself. Thus, gradient descent is only used as a local refinement step, not as a meta-learning procedure. This two‐stage procedure combines cheap random exploration with targeted local search to maximize the prior effectively without incurring the high cost.
Remarks. (i) The specific prior maximization run in parallel; complexity scales linearly with $D$. (ii) Sampling-and-refine (Haar initialization with a few local GD steps) applies per class and per band. (iii) A light regularizer can discourage degenerate reuse of the same orientation across classes, e.g., $\sum_{k}\sum_{d\neq d^{\prime}}\|\mathbf{U}^{\square(d)}_{k}-\mathbf{U}^{\square(d^{\prime})}_{k}\|_{F}^{2}$ when desired.
E.2 Construction Details for Anisotropic Point Cloud Networks
We next consider adaptive canonicalization in the spatial domain for points embedded in ${\mathbb{R}}^{3}$ with xyzxyz coordinates, such as molecular graphs [Duval et al., 2023a] or point clouds [Luo et al., 2022], where each node is associated with a 3D position vector. In this paper, we specifically focus on point cloud networks applied to classification tasks.
Point cloud networks operate directly on xyzxyz coordinates, creating an inherent anisotropy problem [Wiersma et al., 2022, Qian et al., 2021]. The models treat axis-labeled xyzxyz coordinates and their offsets as separate components, with linear layers processing each axis independently. However, under 3D rotation, these coordinate components mix, and since generic weight matrices do not commute with rotation, the model’s response becomes dependent on orientation [Shen et al., 2021].
To address this rotational ambiguity in point cloud classification tasks, we introduce an adaptive canonicalization module that searches over group actions in $\mathcal{SO}(3)$, the 3D rotation group. We now present the application of adaptive canonicalization (AC) to three widely used point cloud learning architectures: DeepSet [Zaheer et al., 2017], PointNet [Qi et al., 2017], and DGCNN [Wang et al., 2019], and denote them respectively AC-DeepSet, AC-PointNet, and AC-DGCNN.

### v2

@@ B.1 Universal Approximation of Euclidean Fucntions
Theorem 14 (Universal Approximation Theorem [LESHNO1993861]).
@@ B.2 Universal Approximation of Multi-Sets Functions
For example, DeeptSets are universal approximators of invariant $C_{0}(\mathcal{K},{\mathbb{R}}^{D})$ functions [wagstaff2022universal].
@@ Appendix C Additional Examples of Continuous Prior Maximization
The above examples can be naturally extended to additional image transformation, like translations and dilations. Notably, translations and dilations do not form a compact group, but still they satisfy the conditions of our theory, which requires no compactness assumptions.
@@ Appendix D Functional Calculus and Spectral Filters
applies the graph convolution theorem [bracewell1966fourier] with $d^{\prime}$ output channel: each Fourier mode $v_{i}$ is preserved in space, while channels are mixed by $g(\lambda_{i})$ in the spectral domain. In the scalar case ($T=d^{\prime}=1$) with $f:\mathbb{R}\to\mathbb{R}$, the spectral filter simply reduces to the functional-calculus operator acting on $\mathbf{X}$:
@@
Spectral graph neural networks [defferrard2017convolutional, kipf2016semi, levie2018cayleynets] compose such filters with pointwise nonlinearities, using trainable $g$ at each layer.
Appendix E Application of Adaptive Canonicalization: Tutorial for Practitioners
@@
In this section, we present the construction details of the application of adaptive canonicalization to anisotropic geometric networks. We start by describing the one vs. rest classifier and our maximization method.
E.1 One vs. Rest Classifiers
@@
In our setting, each output channel $f_{d}\circ\rho_{f_{d}}^{d}(g)\in[0,1]$ is a binary classifier, i.e., representing the probability of $g$ being in class $d$ vs. not being in class $d$ [rifkin2004defense, galar2011overview, allwein2000reducing]. The per-class score is obtained by $\hat{y}_{d}=\sigma(f_{d}\circ\rho_{f_{d}}^{d}(g))$, where $\sigma$ is a sigmoid function. Note that the vector $(\hat{y}_{1},\ldots,\hat{y}_{D})$ is not a probability measure, since each entry represents the independent probability of class $c$ being present. We use binary cross-entropy per class and sum over classes $\sum_{d=1}^{D}(-\,y_{d}\log\hat{y}_{d}-(1-y_{d})\log(1-\hat{y}_{d}))$, where $y_{d}$ denotes the true class.
E.2 Random Maximization
Algorithm 1 Random maximization
Input: Input $g$, backbone network $f$, scalar prior $h(x)$, sampler $\mathtt{Sample}\_\mathtt{U}()$ for $u\sim P$ sampled from a probability measure over $\mathcal{U}$ , number of random samples $K$, gradient descent step $\mathtt{GD}\_\mathtt{Step}$
function $\mathtt{Random}\_\mathtt{Maximization}(g,f,h,\mathtt{Sample}\_\mathtt{U},K,\mathtt{use}\_\mathtt{GD},T,\eta,\mathtt{project}\_\mathtt{to}\_\mathtt{U})$
  $\{u^{1},\ldots,u^{K}\}\leftarrow\mathtt{Sample}\_\mathtt{U}(K)$
  $u^{*}\leftarrow\underset{\{u_{1},\ldots,u_{K}\}}{\arg\max}\;h\circ f(\kappa_{u^{i}}(g))$
  $u^{*}\leftarrow\mathtt{GD}\_\mathtt{Step}(u^{*})$
  return $u^{*}$
We estimate the prior maximization by searching over a transformation space (e.g., unitary orientations per spectral band or rotation for point clouds) and selecting the orientation that maximizes the chosen prior with the classification objective (i.e., one vs. rest). Specifically, we consider a probability measure on the space $\mathcal{U}_{j}$ and draw i.i.d samples $\{u_{j}^{i}\}_{i=1}^{K}$ from it. The argmax of $\{h_{d}\circ f_{d}\big(\kappa_{u_{j}}(g)\big)\}_{u_{j}\in\mathcal{U}_{j}}$ of Eq. (2) is estimated as the argmax of
$\{h_{d}\circ f_{d}\big(\kappa_{u_{j}^{i}}(g)\big)\}_{i=1}^{K}.$
For example, in anisotropic nonlinear spectral filters, we draw a finite pool of candidate transformations from the Haar measure on $\mathcal{U}_{j}$ [mezzadri2006generate]. For anisotropic point cloud networks, the search over rotations can be implemented with quaternion [shoemake1992uniform] or Rodrigues’ formula [rodrigues1840lois]. For each input, we evaluate the prior objective for all candidates in the pool and keep only the maximizing orientation when computing the forward pass and gradients. In this way, the prior maximization is implemented as a randomized search: we sample a set of transformations, apply them in parallel, and pick the one giving the best prior value. Note that this sampling-based maximization is best thought of as one convenient implementation of our framework rather than a requirement, as other optimization strategies over $\mathcal{U}_{j}$ could be implemented and plugged as long as they approximately solve the same maximization problem.
To understand how well this random maximization approximates the ideal maximization over the full transformation space, we can adapt the analysis from cordonnier2024convergence. Their results provide a tail bound for approximating a maximum over a probability space by the maximum over random i.i.d. samples.
We first recall the notion of the volume retaining probability space introduced by cordonnier2024convergence.
Definition 21 (Volume retaining property [cordonnier2024convergence]).
Let $X\subset{\mathbb{R}}^{d}$ and let $P$ be a probability measure on $X$. We say that the probability space $(X,P)$ has the $(r_{0},\kappa)$-volume retaining Lebesgue measure if there exist constants $r_{0}>0$ and $\kappa>0$ such that for any $r\leq r_{0}$ and any $x\in X$
$P(B(x,r)\cap X)\geq\kappa\lambda_{d}(B(x,r)),$
where $\lambda_{d}$ is the $d$-dimensional Lebesgue measure and $B(x,r)$ is the ball center at $x$ with radius $r$.
In our case, points are randomly sampled from some canonical measure over $\mathcal{U}_{j}$ (i.e., the Haar measure), and in all of our example applications $\mathcal{U}_{j}$ has the volume retaining property.
On a volume retaining space, cordonnier2024convergence prove the following concentration inequality for maxima.
Lemma 22 (Concentration inequality for volume retaining space [cordonnier2024convergence]).
Let $(X,P)$ be a probability space with the $(r_{0},\kappa)$-volume retaining property and let $g:X^{2}\rightarrow{\mathbb{R}}^{q}$ be $K_{g}$-Lipschitz. For any $\rho\geq\exp(-n\kappa r_{0}^{d}2^{d})$, for any random variables $X_{1},\ldots,X_{n}\overset{\text{i.i.d.}}{\sim}P$, with probability at least $1-\rho$, it holds
$\|\underset{1\leq i\leq n}{\max}g(x,X_{i})-\sup g(x,\cdot)\|_{\infty}\leq\frac{K_{g}}{2}\left(\frac{\ln(q/\rho)}{n\kappa}\right)^{1/d}.$
Applying this lemma to our setting, $g$ is the output of the task neural network on some class. Since typical neural networks are Lipschitz continuous (e.g., any multilayer perceptron based on a Lipschitz activation function), this immediately gives a guarantee that our random maximization method approximates the true maximum. We plan to extend this analysis for future work.
We note that prior maximization can be strengthened by locally refining the sample points $u_{j}^{i}$. Specifically, we initialize with random transformations drawn from the probability measure, perform the sampling-based prior maximization to select a candidate, and then run a few steps of gradient descent to further decrease the objective locally. If $\mathcal{U}$ is a manifold embedded in Euclidean space, one can apply the vanilla Euclidean gradient descent and then project the obtained transformation back to $\mathcal{U}$, or apply Riemannian gradient descent [absil2008optimization] to stay on $\mathcal{U}$. We summarize the random maximization in Alg. 1. In our experiments, we use 32 sampled transformations for anisotropic nonlinear spectral filters and use 50 sampled transformations for point clouds.
Finally, we note that when the prior maximization has error $e$ (which is a random variable), this leads to an additive term $e$ in universal approximation. Namely, for any $\epsilon>0$, any continuous function can be approximated by $\theta\circ\rho_{\theta}$ up to error $\epsilon+e$ instead of $\epsilon$ in Theorem 6.
E.3 Construction Details for Anisotropic Nonlinear Spectral Filters
In spectral methods for graphs, we often use eigenvectors as a core component for graph representation learning. However, these eigenvectors are not uniquely defined. For each eigenvector we can flip its sign, and when an eigenvalue has multiplicity larger than one, any orthogonal basis of its eigenspace is valid. In practice, this means that the same graph can produce different eigenvectors depending on the eigensolver or numerical details, and a spectral graph neural network that takes these eigenvectors as input may give different outputs for the same graph, which affects stability and can reduce performance. Since the model should depend only on the graph structure and not on arbitrary choices of eigenvector bases, i.e., it should be invariant to the choice of the eigenbasis, it is important to remove such ambiguities. In this work, we focus on graph classification tasks with $D$ classes and study how to resolve eigenbasis ambiguities in this setting.
To study this, we work in the following setup. Consider a graph $G=([N],{\mathbf{A}},{\mathbf{S}})$, where $[N]$ is the set of $N$ vertices, ${\mathbf{A}}\in\mathbb{R}^{N\times N}$ is the adjacency matrix, and ${\mathbf{S}}\in\mathbb{R}^{N\times T}$ is an array of node features (row $n$ us the $T$-dimensional feature at node $n$). We consider the normalized graph Laplacian $\mathcal{L}=\mathbf{I}-\mathbf{D}^{-\frac{1}{2}}\mathbf{A}\mathbf{D}^{-\frac{1}{2}}$ as the graph shift operator (GSO) [^2] Note that any other self-adjoint GSO can be chosen. in our experiments, where $\mathbf{D}$ is a diagonal degree matrix with diagonal entries $(d_{i})_{i}\in\mathbb{R}^{N}$, where $d_{i}$ is the degree of node $i$. The eigendecomposition of $\mathcal{L}$ is given by $\mathcal{L}={\mathbf{V}}^{(G)}\mathbf{\Lambda}{\mathbf{V}}^{(G)\top}$, where ${\mathbf{V}}^{(G)}=({\mathbf{v}}_{i})_{i}\in{\mathbb{R}}^{N\times N}$ is an orthogonal matrix of eigenvectors as the columns (i.e., an eigenbasis) and $\mathbf{\Lambda}=\operatorname{diag}(\lambda_{1},\ldots,\lambda_{N})$ is the diagonal matrix of eigenvalues, where $0\leq\lambda_{1}\leq\ldots\leq\lambda_{N}\leq 2$. We then group the spectrum into predefined bands with boundaries $b_{0}<b_{1}<\cdots<b_{B}$ contained in $[0,2]$, where $B\in\mathbb{N}$ is the total number of bands. The total band $b_{0},b_{B}]$ is a subset od $[0,2]$ since the spectrum of the normalized Laplacian lies in this interval [^3] For other GSOs, different bands can be chosen to match their spectral range, or to cover a subset of the spectrum.. In our implementation, we use a dyadic partitioning scheme. Given a decay rate $0<r<1$, we set
$b_{0}=0,\;b_{k}=2r^{B-k}\text{ for }k=1,\ldots,B-1,\;b_{B}=2,$
and define the $k$-th band as $b_{k-1},b_{k})$. A larger $B$ yields narrower bands and hence finer spectral resolution. For example, taking $B=5$ and $r=0.5$ gives five bands: $[0,0.125)$, $[0.125,0.25)$, $[0.25,0.5)$, $[0.5,1)$ and $[1,2)$.
We now make explicit the symmetry we want the model to respect. Recall that the eigendecomposition of the normalized Laplacian is $\mathcal{L}={\mathbf{V}}^{(G)}\mathbf{\Lambda}{\mathbf{V}}^{(G)\top}$, where ${\mathbf{V}}^{(G)}=({\mathbf{v}}_{i})_{i}\in\mathbb{R}^{N\times N}$ collects the eigenvectors and $\mathbf{\Lambda}=\mathrm{diag}(\lambda_{1},\ldots,\lambda_{N})$ the eigenvalues. Given the band boundaries $b_{0}<\cdots<b_{B}$, we define for each band $k$ the index set $I_{k}(G):=\{i\in[N]:\lambda_{i}\in[b_{k-1},b_{k})\}$, and let ${\mathbf{V}}_{k}^{(G)}\in\mathbb{R}^{N\times M_{k}(G)}$ be the submatrix of ${\mathbf{V}}^{(G)}$ whose columns are the eigenvectors $(\mathbf{v}_{i})_{i\in I_{k}(G)}$. Therefore, we can write ${\mathbf{V}}^{(G)}=[\,{\mathbf{V}}_{1}^{(G)}\,|\,\cdots\,|\,{\mathbf{V}}_{B}^{(G)}]$. For each band we denote the associated Paley-Wiener space by $\mathcal{X}_{k}(G):=\operatorname{span}\{{\mathbf{v}}_{i}:i\in I_{k}(G)\}=\operatorname{Im}({\mathbf{V}}_{k}^{(G)})$. The ambiguity we want to handle comes from changing the orthonormal basis inside each band. Any other basis that of the $k$-th band has the form $\widetilde{{\mathbf{V}}}_{k}^{(G)}={\mathbf{V}}_{k}^{(G)}\,{\mathbf{U}}_{k}^{(G)}$, where ${\mathbf{U}}_{k}^{(G)}\in{\mathbb{R}}^{M_{k}(G)\times M_{k}(G)}$ is a unitary matrix. Equivalently, we can write a full orthonormal basis of $\mathbb{R}^{N}$ whose columns are partitioned into subsequences corresponding to the bands, and each subsequence spans the associated Paley-Wiener space $\mathcal{X}_{k}(G)$. We can write this condition in matrix form as: $\mathbb{R}^{N\times N}\ni\widetilde{{\mathbf{V}}}^{(G)}={\mathbf{V}}^{(G)}{\mathbf{U}}^{(G)}$, where ${\mathbf{U}}^{(G)}:=\operatorname{diag}({\mathbf{U}}_{1}^{(G)},\ldots,{\mathbf{U}}_{B}^{(G)})\in{\mathbb{R}}^{N\times N}$, i.e., the a block matrix with diagonal blocks ${\mathbf{U}}_{k}^{(G)}$. The set of block-diagonal unitary matrices $\mathcal{U}^{(G)}=\{{\mathbf{U}}^{(G)}=\operatorname{diag}({\mathbf{U}}_{1}^{(G)},\ldots,{\mathbf{U}}_{B}^{(G)})\}$ constitutes the band-preserving transformations, i.e., the unitary matrices that keep the Paley-Wiener spaces (or bands) invariant in the sense that ${\mathbf{U}}^{(G)}{\mathbf{v}}\in\mathcal{X}_{k}(G),$ $\forall\,k\in[B],\ \forall\,{\mathbf{v}}\in\mathcal{X}_{k}(G)$. The space of these matrices constitute the symmetry space in our setting.
The above construction is defined for a single graph, but in our setting we work with a collection of graphs. The spectral support of the bands, that is, the intervals $b_{k-1},b_{k})$, is fixed across all graphs. However, for a given graph $G$, the $k$-th band, i.e, the linear span of the eigenvectors whose eigenvalues lie in $b_{k-1},b_{k})$, and this subspace depends on $G$. Different graphs can have a different number of eigenvalues in the same interval, so the dimension of the $k$-th band is not constant across graphs. We denote by $M_{k}(G)$ the dimension of the $k$-th band for the graph $G$.
In order to apply a task network given by a standard neural network architecture, such as an MLP, we require that all inputs lie in a common vector space of fixed dimension, independent of the particular graph $G$. The variability of the band dimensions $M_{k}(G)$ across graphs violates this requirement. To resolve this, we introduce a padding operator that standardizes the size of each band. Implementation-wise, for each band $k$ we fix for each band $k$ the size to be the largest value of $M_{k}(G)$ over the dataset and denote this value by $J_{k}$, and we zero-pad bands that are smaller than this size. Let $J_{1},\ldots,J_{B}$ be these integers, and denote $J=\sum_{k=1}^{B}J_{k}$. For a graph $G$ with signal $\mathbf{S}\in{\mathbb{R}}^{N\times T}$, let $C_{k}(G,\mathbf{S})\in\mathbb{R}^{M_{k}(G)\times T}$ be the matrix of spectral coefficients in the $k$-th band (one row per eigenvector in the band and $T$ channels). Since $M_{k}(G)\leq J_{k}$ by construction, we define the band-wise padding operator $P_{k}$ by
$P_{k}(G,\mathbf{S})\coloneqq\begin{bmatrix}C_{k}(G,\mathbf{S})\\[1.99997pt] \mathbf{0}_{(J_{k}-M_{k}(G))\times T}\end{bmatrix}\in\mathbb{R}^{J_{k}\times T}.$
Collecting all bands, we obtain the global operator $P(G,\mathbf{S})\coloneqq(P_{k}(G,\mathbf{S}))_{k=1}^{B}\in\mathbb{R}^{J\times T}$. By construction, $P$ maps every graph–signal pair $(G,\mathbf{S})$ to a spectral representation in the fixed-dimensional vector space $\mathbb{R}^{J\times T}$, so that we can apply the same task network $\Psi:\mathbb{R}^{J\times T}\to\mathbb{R}^{D}$ (e.g., an MLP) uniformly across all graphs. After standardizing the band sizes to $J_{k}$ and writing $J=\sum_{k=1}^{B}J_{k}$, we use the same parameterization with $\{{\mathbf{U}}=\operatorname{diag}({\mathbf{U}}_{1},\ldots,{\mathbf{U}}_{B})\in{\mathbb{R}}^{J\times J},\text{ where }{\mathbf{U}}_{k}\in{\mathbb{R}}^{J_{k}\times J_{k}}\text{ is unitary}\}$, which gives a parameterization of the space of unitary operators that keep the bands invariant.
To resolve the band-wise eigenbasis ambiguity, we introduce the anisotropic nonlinear spectral filters as follows. Consider a symmetryless neural network $\Psi:{\mathbb{R}}^{J\times T}\rightarrow{\mathbb{R}}^{D}$ that operates on the space of spectral coefficients of the signal. We denote $\Psi=(\Psi_{d})_{d=1}^{D}$ where $\Psi_{d}:{\mathbb{R}}^{J\times T}\rightarrow{\mathbb{R}}$. The prior maximization for each class $d\in\{1,\dots,D\}$ is performed by
$\{\mathbf{U}^{\square(d)}_{1},\ldots,\mathbf{U}^{\square(d)}_{B}\}=\underset{\mathbf{U}^{(d)}_{1},\ldots,\mathbf{U}^{(d)}_{B}}{\arg\max}\;h_{d}(\Psi_{d}(\mathbf{U}^{(d)}_{1}(G,\mathbf{S})),\ldots,\mathbf{U}^{(d)}_{B}(P_{B}(G,\mathbf{S})))),$
where $\mathbf{U}^{(d)}_{k}\in\mathbb{R}^{J_{k}\times J_{k}}$ is a unitary matrix and $h_{d}(x)=x$ is class-$d$ prior. Once we obtain the set of the optimal unitary matrices $\{\mathbf{U}^{\square(d)}_{1},\ldots,\mathbf{U}^{\square(d)}_{B}\}$, the class $d$ score is then computed by $\Psi_{d}(\mathbf{U}^{\square(d)}_{1}(P_{1}(\mathbf{V}^{(G)},\mathbf{S})),\ldots,\mathbf{U}^{\square(d)}_{B}(P_{B}(\mathbf{V}^{(G)},\mathbf{S})))$ are then passed through a sigmoid nonlinearity to obtain class probabilities. The training loss is the sum of $D$ binary cross-entropies.
E.4 Construction Details for Anisotropic Point Cloud Networks
We next apply adaptive canonicalization in the spatial domain, where inputs are multisets of points in ${\mathbb{R}}^{3}$ with $(x,y,z)$ coordinates. Specifically, we focus on multiset networks applied to classification tasks. Standard multiset networks, e.g., DeepSet [zaheer2017deep] and PointNet [qi2017pointnet], which take the 3D coordinates of the points as their features, suffer from an inherent anisotropy problem [wiersma2022deltaconv, qian2021assanet]: the classification outcomes for a point cloud and for a rotated version of the same point cloud do not coincide in general. To address this sensitivity to orientation in multiset networks, we introduce an adaptive canonicalization module that searches over the 3D rotation group $\mathcal{SO}(3)$.
Let ${\mathbf{X}}\in{\mathbb{R}}^{N\times 3}$ represent a point cloud of $N$ points in $\mathbb{R}^{3}$ and let $\Psi:{\mathbb{R}}^{N\times 3}\rightarrow{\mathbb{R}}^{D}$ be a multiset neural network. In classification, $D$ is the number of classes. We denote $\Psi=(\Psi_{d})_{d=1}^{D}$ where $\Psi_{d}:{\mathbb{R}}^{N\times 3}\rightarrow{\mathbb{R}}$. Here, $\Psi$ is invariant to permutations of the $N$ points, but not invariant to 3D rotations of the point cloud. The prior maximization for each class $d\in\{1,\dots,D\}$ is performed by
${\mathbf{R}}_{d}^{\square}=\underset{{\mathbf{R}}\in\mathcal{SO}(3)}{\arg\max}\;h_{d}(\Psi_{d}({\mathbf{X}}{\mathbf{R}}^{\top})),$
where class-$d$ prior is $h_{d}(x)=x$. Once we found the canonical form ${\mathbf{R}}_{d}^{\square}$, the class scores in our anisotropic point cloud network are computed as $s_{d}=\Psi_{d}({\mathbf{X}}{\mathbf{R}}_{d}^{\square\top})$ and are then passed through a sigmoid nonlinearity to obtain class probabilities. The training loss is the sum of $D$ binary cross-entropies.
The canonicalization setting is independent of the specific multiset neural network architecture. Below, we present three widely used multiset neural networks equipped with adaptive canonicalization.